# Edge Formation Mechanism of Single-Walled Carbon Nanotube Revealed by Molecular Dynamics Simulations based on a Neural Network Potential

### v3

@@ Abstract
Despite the high potential for applications utilizing the unique properties of single-walled carbon nanotubes (SWCNTs), the applications remain challenging due to the difficulty of synthesizing SWCNTs with a specific chirality. To elucidate the mechanisms that determine their chirality during growth, intensive efforts have been devoted to classical molecular dynamics (MD) simulations. However, the mechanism of chirality determination has not been fully clarified, which can partly be attributed to the limited accuracy of empirical interatomic potentials in reproducing the behavior of carbon and metal atoms. In this work, we develop a neural network potential (NNP) for carbon-metal system to accurately describe the SWCNT growth, and perform MD simulations of SWCNT growth using the NNP. The MD simulations demonstrate the process of defect-free chirality-definable SWCNT growth with dynamic rearrangement of edge configurations, and the consistency between the probability of edge configuration appearance and the entropy-driven edge stability predicted from the nanotube-catalyst interfacial energy. It is also shown that the edge defect formation is induced by vacancy and suppressed by vacancy healing through adatom diffusion on the SWCNT edges. These results provide insights into the edge formation thermodynamics and kinetics of SWCNTs, an important clue to the chirality-controlled synthesis of SWCNTs.
@@ 1 Introduction
Single-walled carbon nanotubes (SWCNTs)1 have great potential as a next-generation material for nanoscale devices such as sensors, optoelectronic devices, and high-performance logic transistors2, 3 because of their excellent physical properties4. The electronic structure of SWCNTs dramatically varies from metallic to semiconducting depending on their chirality5, 6, 7. chirality-controlled synthesis is important to obtain chirality-pure SWCNT samples without contamination and damage caused by sorting, and to realize high-end SWCNT-based devices. Many experimental studies have been carried out to achieve chirality-controlled growth of SWCNTs8. Although these impressive results shed light on chirality control, further understanding of the chirality-determining SWCNT growth mechanism is needed to produce higher purity SWCNT samples with arbitrary chirality in large quantity.
@@
In the past few decades, theoretical models have been proposed to account for the chirality distribution of SWCNTs9, 10. These models successfully relate the state of the interface between the SWCNT edge and the catalyst to the chirality distribution, but do not provide a detailed atomistic mechanism in the edge formation process because they are based on some assumptions about the process of carbon addition to the SWCNT edges. To reveal such mechanism, many molecular dynamics (MD) simulations have also been performed. In particular, classical MD simulations, which are suitable for observing nanotube growth on a long time scale, have demonstrated a series of growth processes from the nucleation to the sidewall elongation11, 12, 13, 14, 15, 16, 17. As a first approach to the discussion of chirality in MD simulation, Neyts et al.18, 19 performed hybrid atomistic simulations of MD and force-biased Monte Carlo based on the reactive force field (ReaxFF). They defined the local chirality by the diameter and the angle of the hexagon to the axial direction of SWCNTs. In their simulations of the initial growth process of SWCNTs, chirality changes from zigzag to armchair were observed. Yoshikawa et al.20 carried out MD simulations with a lower carbon supply rate, and assigned the chirality to SWCNTs by defect-free growth using the Tersoff-type potential. However, the chiralities of the observed SWCNTs were only zigzag and near-zigzag, which is contrary to the experimental reports of near-armchair preferential growth21, 22, 23, 24. This suggested that the bond-order potential shows an excessive preference for zigzag edges, as reported in a theoretical study25, and that the preference for armchair edges over zigzag edges observed in the ab initio calculation cannot be correctly reproduced in the bond-order potential.
@@
Currently, there are two possible reasons why it is difficult to simulate the growth of near-armchair SWCNTs while maintaining their chirality without defects, as reported in the experimental study26. Firstly, the time scale is too short for defect healing. Secondly, the energetics of SWCNTs including defects is not accurately described by classical interatomic potentials. In order to solve these problems and to reproduce defect-free growth, it is necessary to perform MD simulations on a long time scale using an accurate interatomic potential.
@@
Recently, the application of artificial neural networks to interatomic potentials has been actively studied to overcome the limitations of conventional interatomic potentials and such interatomic potentials are called neural network potentials (NNPs). Since the idea of symmetry functions to represent high-dimensional potential energy surfaces (PESs) with translational and rotational invariance was proposed by Behler and Parrinello27, a number of approaches have been developed to construct NNPs in the field of materials science28. Among the NNPs, the deep potential29 is one of the most popular methods because of its high accuracy and high computational efficiency.
@@
Here we develop an NNP based on the deep potential for C-Fe system to accurately simulate the SWCNT growth. The NNP is trained against iteratively collected 44000 structures, achieving, among other feartures, good accuracy in the defect formation energy of graphene and the thermodynamic stability of C-Fe nanoparticles. Using the developed NNP, we perform MD simulations of the defect-free SWCNT growth by supplying carbon atoms to Fe nanoparticles. The results show how the chirality of SWCNTs and the interfacial energy between the SWCNT and the Fe nanoparticles affect the structural evolution of the nanotube edges. We also discuss the mechanism of edge defect formation and healing, which leads to the defect-free growth.
@@ 2.1 Validation of NNP for Fe-catalyzed SWCNT Growth
The growth process of SWCNTs consists of the dynamic formation of carbon rings on the catalyst surface through recombining inter-carbon bonds. In order to reproduce the SWCNT growth by MD simulation, interatomic potentials need to predict properties of various morphology of C-Fe system including unstable structures. In this study, we used classical MD simulation for sampling dataset structures and trained an NNP that can accurately predict the energies and forces of various structures including molecules, chains, clusters, graphene, and nanotubes. Before focusing on the mechanism of SWCNT growth, we validated the NNP from the perspectives of defect formation energy and edge free energy of graphene, thermodynamic behavior of C-Fe nanoparticles.
@@ 2.1.1 Defect formation energies of graphene
Here, $E_{d}$ and $E_{g}$ are the total energy of defective and ideal graphene structure, respectively. $n_{d}$ and $n_{g}$ are the number of atoms in the supercell of the defective and ideal graphene, respectively. The comparison of defect formation energies between density functional theory (DFT) calculation and the NNP is plotted in Figure 1a. The same plots for Tersoff-type potential30 and ReaxFF31, 32, employed in previous SWCNT growth simulations, are also shown. The plots indicate that the developed NNP more accurately predicted the defect formation energy than the other two interatomic potentials. The Tersoff-type potential underestimated the defect formation energy, which is considered to result in more frequent defect formation during SWCNT growth simulations30, 20. In contrast, the ReaxFF potential overestimated the defect formation energy and accordingly the stability of graphene, which may have prevented cap lift-off driven by pentagon formation during the initial stages of SWCNT growth.
@@ 2.1.2 Edge energy of graphene
The edge energy is considered to be an important factor of SWCNT chirality selectivity25, 33. To verify the NNP and discuss the chirality selectivity in terms of thermodynamic stability, we checked the chiral angle dependence of the edge energy of graphene. To calculate the edge energy by the NNP, we modeled graphene nanoribbons (GNRs) in the cell with one periodic direction and a vacuum space of 15 Å in the other two non-periodic directions. We relaxed them with the conjugate gradient method. The edge energy $\gamma$ is defined as
@@
where $E_{total}$ is the total energy of the supercell, $N_{C}$ is the total number of carbon atoms in the supercell, $E_{C}$ is the energy of carbon atoms in the graphene lattice, $N_{e}$ is the number of terminating atoms, $\mu$ is the chemical potential of the terminating element, and $L$ is the periodic length of graphene nanoribbon as defined in a previous theoretical study25. We computed the edge energy of the GNRs plotted as a function of chiral angle, as shown in Figure 1b. For comparison, we also plotted the analytical curve 25 expressed as
@@
where $\chi$ is the chiral angle, $\gamma^{\prime}_{A}$ and $\gamma^{\prime}_{Z}$ are the edge energies of armchair and zigzag edge, respectively. On pristine graphene edges, the energies computed with the NNP showed higher stability of large chiral angles, which agreed with the generalized gradient approximation (GGA) of DFT. On the other hand, those computed with the Tersoff-type potential indicated higher stability of small chiral angles, which was considered to result in the zigzag preferential growth by previous MD simulations20.
Figure 1: Comparison of (a) defect formation energy of graphene and (b) edge energy of graphene as a function of chiral angle between DFT with GGA, the developed NNP, Tersoff-type potential30, and ReaxFF31, 32. The edge energies of graphene were directly computed (dots) and fitted with Eq. 3 (lines).
@@ 2.1.3 Thermodynamic properties of Fe
The influence of the catalyst phase on SWCNT growth has been reported in experimental studies34, 35. In order to accurately simulate SWCNT growth, the thermodynamic behavior of catalyst nanoparticles must be well considered. This section discusses the thermodynamic behavior of Fe bulk crystals and nanoparticles. First, we estimated the melting point of bulk body centered cubic (bcc) Fe by two-phase simulation36 using the NNP. For two-phase simulation, we prepared the initial structure of bulk Fe with a bcc-liquid interface by melting the upper half of a bcc supercell of $28.52\times 28.52\times 74.17$ with 5200 atoms at 5000 K (Figure S3a). Starting with the initial structures, we carried out MD simulations for 200 ps under NPT ensemble at 1 atm with different temperatures. The final structures at 1850, 1887.5, and 1900 K are shown in Figure S3b-d, and the potential energy was well converged at each temperature, as shown in Figure S3e. The equilibrium temperature was 1887.5$\pm$12.5 K, which agrees well with the experimental melting point of 1811 K37. In addition, the latent heat of melting was estimated as 0.15 eV/atom from the difference of the averaged potential energies between solid and liquid, which also agrees with the experimental value 0.143 eV/atom38 (Figure S3f).
@@
where $r_{ij}$ is the distance between atom $i$ and $j$, N is the number of particles and the bracket is the time average calculated over MD simulation at constant temperature $T$. The melting point was determined as the temperature at which the Lindemann index changed discontinuously. Previous studies used a threshold value in the range of 0.08-0.15 as the melting point criteria39, 40, 41, 42, 43, 44. However, in the case of small nanoparticles, the solid and liquid phases coexist over a wide temperature range, and the criteria for the melting point have not been well established. In this study, to estimate the melting point, a hyperbolic function was fitted to the data point of the Lindemann index. The mathematical form of the hyperbolic function is expressed as
@@
where $a$, $b$, $c$, $d$, $e$, $s$, $v$, and $w$ are fitting parameters. The melting point was determined as $T_{m}=w$ where the inflection point of the fitted hyperbolic curve is located (see Figure S4, Supporting Information). For MD simulation to obtain the Lindemann index, the initial structures were prepared by melting Fe nanoparticles at 2000 K and then annealing them to 0 K at 1 K/ps using the NNP. The annealed structures of Fe200, Fe600 and Fe2000 are shown in Figures 2a-c. The stable phase of Fe nanoparticles changed depending on the size. For large nanoparticles ($N>$700), the bcc structure was the stable phase as well as bulk Fe. However, for medium-size nanoparticles (300$<N<$700), icosahedral (Ih) was found stable. No regular structures were not observed for smaller nanoparticles ($N<$300) as not all possible magic number structures were used as a starting point in this first approach. This phase transition from bcc to Ih by size, which was considered to be due to surface effect, is consistent with high-resolution transmission electron microscopy observation of Ih Fe nanoparticles45. Starting from the annealed structures, MD simulations were performed for 2 ns at constant temperature $T$ and the Lindemann index was then obtained as the average over the latter half 1 ns of MD simulations. According to the Gibbs-Thomson equation, the decrease in the melting point is inversely proportional to the diameter. As shown in Figure 2d, the melting point followed the equation for large nanoparticles of bcc Fe. However, the melting point discontinuously changed by 300 K between Fe600 and Fe800. This discontinuity of the melting point corresponded to the phase transition from bcc to Ih.
In addition, the melting points of Fe120Cn were calculated, as shown in Figure 2e. The melting point decreased from 814 K to 740 K with $\sim$2.5 % carbon content, while it increased with more carbon content. The observed V-shape diagram is consistent with the phase diagram of bulk Fe-C46 and previous computational studies39, 47.
Figure 2: Thermodynamic behaviors of Fe nanoparticles. (a-c) Snapshots of cross sectional views of annealed structures of Fe200, Fe600 and Fe2000 nanoparticles. The blue and red balls denote atoms in bcc and Ih structure, respectively. The white balls denote atoms out of bcc and Ih structures. The structures were categorized by the common neighbor analysis48. (d) Melting point of Fe nanoparticles as a function of inverse diameter. The red line is linear fits to data points for bcc Fe nanoparticles. (e) Melting points of Fe120Cn nanoparticles obtained by adding up to 17.5 % of carbon atoms.

### v5

# Edge Dynamics in Iron-Cluster Catalyzed Growth of Single-Walled Carbon Nanotubes Revealed by Molecular Dynamics Simulations based on a Neural Network Potential

@@ Abstract
Given the high potential for applications utilizing the unique properties of single-walled carbon nanotubes (SWCNTs), there is considerable enthusiasm for addressing the challenges associated with synthesizing SWCNTs with specific chirality. To elucidate the mechanisms that determine the chirality of SWCNTs during growth, intensive efforts have been devoted to classical molecular dynamics (MD) simulations. However, the mechanism of chirality determination has not been fully clarified, which can partly be attributed to the limited accuracy of empirical interatomic potentials in reproducing the behavior of carbon and metal atoms. In this work, we develop a neural network potential (NNP) for carbon-metal system to accurately describe the SWCNT growth, and perform MD simulations of SWCNT growth using the NNP. The MD simulations illustrate the defect-free, chirality-definable growth of SWCNTs, highlighting the dynamic rearrangement of edge configurations and the consistency between the probability of edge configuration appearance and the entropy-driven edge stability model proposed here. It is also shown that the edge defect formation is induced by vacancy and suppressed by vacancy healing through adatom diffusion on the SWCNT edges. These results provide insights into the edge formation thermodynamics and kinetics of SWCNTs, an important clue to the chirality-controlled synthesis of SWCNTs.
@@ 1 Introduction
Single-walled carbon nanotubes (SWCNTs) have great potential 1, 2 as a next-generation material for nanoscale devices such as sensors, optoelectronic devices, and high-performance logic transistors3, 4, 5 because of their excellent physical properties4, 6. The electronic structure of SWCNTs dramatically varies from metallic to semiconducting depending on their chirality7, 8, 9, and chirality-controlled synthesis is important to obtain chirality-pure SWCNT samples without contamination and damage caused by sorting, and to realize high-end SWCNT-based devices. Many experimental studies have been carried out to achieve chirality-controlled growth of SWCNTs2. Although these impressive results shed light on chirality control, further understanding of the chirality-determining SWCNT growth mechanism is needed to produce higher purity SWCNT samples with arbitrary chirality in large quantity.
@@
In the past few decades, theoretical models have been proposed to account for the chirality distribution of SWCNTs10, 11. These models successfully relate the state of the interface between the SWCNT edge and the catalyst to the chirality distribution, but do not provide a detailed atomistic mechanism in the edge formation process because they are based on some assumptions about the process of carbon addition to the SWCNT edges. To reveal such mechanism, many molecular dynamics (MD) simulations have also been performed. In particular, classical MD simulations, which are suitable for observing nanotube growth on a long time scale, have demonstrated a series of growth processes from the nucleation to the sidewall elongation12, 13, 14, 15, 16, 17, 18. As a first approach to the discussion of chirality in MD simulation, Neyts et al.19, 20 performed hybrid atomistic simulations of MD and force-biased Monte Carlo based on the reactive force field (ReaxFF). They defined the local chirality by the diameter and the angle of the hexagon to the axial direction of SWCNTs. In their simulations of the initial growth process of SWCNTs, chirality changes from zigzag to armchair were observed. Yoshikawa et al.21 carried out MD simulations with a lower carbon supply rate, and assigned the chirality to SWCNTs by defect-free growth using the Tersoff-type bond-order potential. This chirality-definable MD simulation demonstrated the kink-running process at SWCNT edges predicted by the dislocation theory of SWCNT growth 10. However, the chiralities of the observed SWCNTs were only zigzag and near-zigzag, which is contrary to the experimental reports of near-armchair preferential growth22, 23, 24, 25. This suggests that the bond-order potential shows an excessive preference for zigzag edges, as reported in a theoretical study26, and that the preference for armchair edges over zigzag edges predicted by the DFT calculation is not correctly reproduced by the bond-order potential.
@@
Currently, there are two reasons why it is difficult to accurately simulate SWCNT growth. Firstly, the time scale of MD simulations is smaller than that of experiments, and there is not enough time for defects to heal during simulated SWCNT growth. Secondly, the energetics of SWCNTs including defects is not accurately described by classical interatomic potentials. For more accurate description of atomistic behavior in SWCNT growth, it is necessary to perform long MD simulations using an accurate interatomic potential.
@@
Recently, the application of artificial neural networks to interatomic potentials has been actively studied to overcome the limitations of conventional interatomic potentials and such interatomic potentials are called neural network potentials (NNPs). Since the idea of symmetry functions to represent high-dimensional potential energy surfaces (PESs) with translational and rotational invariance was proposed by Behler and Parrinello27, a number of approaches have been developed to construct NNPs in the field of materials science28. Among the NNPs, the deep potential29 is one of the most popular methods because of its high accuracy and high computational efficiency. Hedman et al. have performed MD simulations of SWCNT growth using the deep potential at the same time, with a different training dataset, independently of this work 30, and healing of interface defects was extensively studied.
@@
In this study, we develop an NNP based on the deep potential for C-Fe system to accurately simulate the SWCNT growth. The NNP is trained against iteratively collected 44000 structures, achieving good accuracy in the defect formation energy of graphene and the thermodynamic stability of C-Fe nanoparticles. Using the developed NNP, we perform MD simulations of the defect-free SWCNT growth by supplying carbon atoms to Fe nanoparticles. The MD simulations demonstrate the process of defect-free chirality-definable SWCNT growth with dynamic rearrangement of edge configurations, and the consistency between the probability of edge configuration appearance and the entropy-driven edge stability predicted from the nanotube-catalyst interfacial energy. We also discuss the mechanism of edge defect formation and healing, which leads to the defect-free growth.
@@ 2.1 Validation of NNP for Fe-catalyzed SWCNT Growth
The growth process of SWCNTs consists of the dynamic formation of carbon rings on the catalyst surface through recombining carbon-carbon bonds. In order to reproduce the SWCNT growth by MD simulation, interatomic potentials need to predict properties of various morphology of C-Fe system including unstable structures. In this study, we used classical MD simulation for sampling dataset structures and trained an NNP that can accurately predict the energies and forces of various structures including molecules, chains, clusters, graphene, and nanotubes. Before focusing on the mechanism of SWCNT growth, we validated the NNP from the perspectives of defect formation energy and edge free energy of graphene, thermodynamic behavior of C-Fe nanoparticles.
@@ 2.1.1 Defect formation energies of graphene
Here, $E_{d}$ and $E_{g}$ are the total energy of defective and ideal graphene structure, respectively. $n_{d}$ and $n_{g}$ are the number of atoms in the supercell of the defective and ideal graphene, respectively. The comparison of defect formation energies between density functional theory (DFT) calculation and the NNP is plotted in Figure 1a. The same plots for Tersoff-type potential31 and ReaxFF32, 33, employed in previous SWCNT growth simulations, are also shown. The plots indicate that the developed NNP much more accurately predicted the defect formation energy than the other two interatomic potentials. The Tersoff-type potential underestimated the defect formation energy, which is considered to result in more frequent defect formation during SWCNT growth simulations31, 21. In contrast, the ReaxFF potential overestimated the defect formation energy and accordingly the stability of graphene, which may have prevented cap lift-off driven by pentagon formation during the initial stages of SWCNT growth.
@@ 2.1.2 Edge energy of graphene
The edge energy is considered to be an important factor of SWCNT chirality selectivity26, 34. To verify the NNP and discuss the chirality selectivity in terms of thermodynamic stability, we checked the chiral angle dependence of the edge energy of graphene. To calculate the edge energy by the NNP, we modeled graphene nanoribbons (GNRs) in the cell with one periodic direction and a vacuum space of 15 Å in the other two non-periodic directions. We relaxed them with the conjugate gradient method. The edge energy $\gamma$ is defined as
@@
where $E_{total}$ is the total energy of the supercell, $N_{C}$ is the total number of carbon atoms in the supercell, $E_{C}$ is the energy of carbon atoms in the graphene lattice, $N_{e}$ is the number of terminating atoms, $\mu$ is the chemical potential of the terminating element, and $L$ is the periodic length of graphene nanoribbon as defined in a previous theoretical study26. We computed the edge energy of the GNRs plotted as a function of chiral angle, as shown in Figure 1b. For comparison, we also plotted the analytical curve 26 expressed as
@@
where $\chi$ is the chiral angle, $\gamma^{\prime}_{A}$ and $\gamma^{\prime}_{Z}$ are the edge energies of armchair and zigzag edge, respectively. On pristine graphene edges, the energies computed with the NNP showed higher stability of large chiral angles, which agreed with the generalized gradient approximation (GGA) of DFT. On the other hand, those computed with the Tersoff-type potential indicated higher stability of small chiral angles, which was considered to result in the zigzag preferential growth by previous MD simulations21.
Figure 1: Comparison of (a) defect formation energy of graphene and (b) edge energy of graphene as a function of chiral angle between DFT with GGA, the developed NNP, Tersoff-type potential31, and ReaxFF32, 33. The edge energies of graphene were directly computed (dots) and fitted with Eq. 3 (lines).
@@ 2.1.3 Thermodynamic properties of Fe
The influence of the catalyst phase on SWCNT growth has been reported in experimental studies35, 36. In order to accurately simulate SWCNT growth, the thermodynamic behavior of catalyst nanoparticles must be well considered. This section discusses the thermodynamic behavior of Fe bulk crystals and nanoparticles. First, we estimated the melting point of bulk body centered cubic (bcc) Fe by two-phase simulation37 using the NNP. For two-phase simulation, we prepared the initial structure of bulk Fe with a bcc-liquid interface by melting the upper half of a bcc supercell of $28.52\times 28.52\times 74.17$ with 5200 atoms at 5000 K (Figure S3a). Starting with the initial structures, we carried out MD simulations for 200 ps under NPT ensemble at 1 atm with different temperatures. The final structures at 1850, 1887.5, and 1900 K are shown in Figure S3b-d, and the potential energy was well converged at each temperature, as shown in Figure S3e. The equilibrium temperature was 1887.5$\pm$12.5 K, which agrees well with the experimental melting point of 1811 K38. In addition, the latent heat of melting was estimated as 0.15 eV/atom from the difference of the averaged potential energies between solid and liquid, which also agrees with the experimental value 0.143 eV/atom39 (Figure S3f).
@@
where $r_{ij}$ is the distance between atom $i$ and $j$, N is the number of particles and the bracket is the time average calculated over MD simulation at constant temperature $T$. The melting point was determined as the temperature at which the Lindemann index changed discontinuously. Previous studies used a threshold value in the range of 0.08-0.15 as the melting point criteria40, 41, 42, 43, 44, 45. However, in the case of small nanoparticles, the solid and liquid phases coexist over a wide temperature range, and the criteria for the melting point have not been well established. In this study, to estimate the melting point, a hyperbolic function was fitted to the data point of the Lindemann index. The mathematical form of the hyperbolic function is expressed as
@@
where $a$, $b$, $c$, $d$, $e$, $s$, $v$, and $w$ are fitting parameters. The melting point was determined as $T_{m}=w$ where the inflection point of the fitted hyperbolic curve is located (see Figure S4, Supporting Information). For MD simulation to obtain the Lindemann index, the initial structures were prepared by melting Fe nanoparticles at 2000 K and then annealing them to 0 K at 1 K/ps using the NNP. The annealed structures of Fe200, Fe600 and Fe2000 are shown in Figures 2a-c. The stable phase of Fe nanoparticles changed depending on the size. For large nanoparticles ($N>$700), the bcc structure was the stable phase as well as bulk Fe. However, for medium-size nanoparticles (300$<N<$700), icosahedral (Ih) was found stable. No regular structures were not observed for smaller nanoparticles ($N<$300) as not all possible magic number structures were used as a starting point in this first approach. This phase transition from bcc to Ih by size, which was considered to be due to surface effect, is consistent with high-resolution transmission electron microscopy observation of Ih Fe nanoparticles46. Starting from the annealed structures, MD simulations were performed for 2 ns at constant temperature $T$ and the Lindemann index was then obtained as the average over the latter half 1 ns of MD simulations. According to the Gibbs-Thomson equation, the decrease in the melting point is inversely proportional to the diameter. As shown in Figure 2d, the melting point followed the equation for large nanoparticles of bcc Fe. However, the melting point discontinuously changed by 300 K between Fe600 and Fe800. This discontinuity of the melting point corresponded to the phase transition from bcc to Ih.
In addition, the melting points of Fe120Cn were calculated, as shown in Figure 2e. The melting point decreased from 814 K to 740 K with $\sim$2.5 % carbon content, while it increased with more carbon content. The observed V-shape diagram is consistent with the phase diagram of bulk Fe-C47 and previous computational studies40, 48. The following simulations of SWCNT growth are performed under conditions where the Fe catalysts are fully liquid.
Figure 2: Thermodynamic behaviors of Fe nanoparticles. (a-c) Snapshots of cross sectional views of annealed structures of Fe200, Fe600 and Fe2000 nanoparticles. The blue and red balls denote atoms in bcc and Ih structure, respectively. The white balls denote atoms out of bcc and Ih structures. The structures were categorized by the common neighbor analysis49. (d) Melting point of Fe nanoparticles as a function of inverse diameter. The red line is linear fits to data points for bcc Fe nanoparticles. (e) Melting points of Fe120Cn nanoparticles obtained by adding up to 17.5 % of carbon atoms.